# Exponential Dichotomies in Higher-Dimensional Spatial Dynamics for Elliptic Partial Differential Equations

### v1

@@ Abstract
Exponential dichotomies, when they exist, provide powerful information about the structure of bounded solutions even in the case of an ill-posed evolutionary equation. The method of spatial dynamics, in which one views a spatial variable as a time-like evolutionary variable, allows for the use of classical dynamical systems techniques, such as exponential dichotomies, in broader contexts. This has been utilized to study stationary, traveling wave, time-periodic, and spiral wave solutions of PDEs on spatial domains with a distinguished unbounded direction (e.g. the real line or a channel of the form $\mathbb{R}\times\Omega$). Recent work has shown how to extend the spatial dynamics framework to elliptic PDEs posed on general multi-dimensional spatial domains. In this paper, we show that, in the same context, exponential dichotomies exist, thus allowing for their use in future analyses of coherent structures, such as spatial patterns in reaction-diffusion equations on more general domains.
Acknowledgements: This material is based upon work supported by the National Science Foundation under Award Numbers DMS-2205434, DMS-2006887, and DMS-2307650.
@@ 1 Introduction
A current challenge in the study of pattern formation is the intractability of analytic dynamical systems approaches for pattern forming partial differential equations in higher spatial dimensions. Dynamical systems techniques, for example center manifold reductions, geometric singular perturbation theory for fast-slow systems, and exponential dichotomies, have been applied to study existence and stability of stationary spatial patterns and traveling wave solutions in a variety of one-dimensional contexts (e.g. [4, 13, 16]).
@@
Spatial dynamics, in its simplest iteration, is the trick of rewriting a second order differential equation into a system of first order equations. For example, $u_{xx}+F(u)=0$ becomes the system
$\begin{array}[]{l}u_{x}=v\\ v_{x}=-F(u).\end{array}$
For PDEs posed on $\mathbb{R}$, this perspective allows stationary and traveling-wave problems to be studied as an ODE, permitting the use of well-understood and widely applicable dynamical system tools. Similarly, PDEs posed on $\mathbb{R}^{n}$ that are restricted to radially symmetric solutions depend only on the radius $r\in\mathbb{R}^{+}$ and can therefore be written as an ODE using spatial dynamics. For example, $\Delta u+F(u)=0$, when written in spherical coordinates with $u=u(r)$, becomes
$\quad u_{rr}+\left(\frac{n-1}{r}\right)u_{r}+F(u)=0$
which is then written as a first order system in the variable $r$ as
$\begin{array}[]{l}u_{r}=v\\ v_{r}=-\left(\frac{n-1}{r}\right)v-F(u).\end{array}$
The nonautonomous $r$-dependence of this radial equation and the singularity that occurs at $r=0$ make this setting meaningfully different from the one-dimensional framework. Scheel made significant contributions to characterizing the behavior of radially symmetric spatial patterns in [19] where he developed a systematic bifurcation theory including extensions of center manifold reduction and normal form theory. In [19], Scheel also established existence of both stationary and time-periodic radially symmetric patterns using these techniques.
When spatial dynamics is applied to elliptic a PDEs posed on a cylindrical domain $\mathbb{R}\times\Omega\subset\mathbb{R}^{n}$, the result is an evolutionary equation posed on an infinite-dimensional function space (rather than a finite dimensional vector space as in the case of ODEs). The same reaction diffusion problem $\Delta u+F(u)=0$ is decomposed into the variables $(x,y)\in\mathbb{R}\times\Omega$ to become
$u_{xx}+\Delta_{\Omega}u+F(u)=0,\qquad u(x,\cdot)\in L^{2}(\Omega).$
The spatial evolutionary equation that results from writing the above as a first-order system in $x$ is
$\begin{array}[]{l}u_{x}=v\\ v_{x}=-\Delta_{\Omega}u-F(u).\end{array}$
There is a long history of extending dynamical systems tools into this infinite-dimensional context (e.g. [6, 7, 8, 11]). These extensions have made possible existence proofs in the channel setting [5], construction of local and global manifolds [14], alternative methods to understand the Fredholm properties of elliptic operators [15], and stability analyses of traveling waves [13].
@@
A similarly infinite-dimensional evolutionary equation emerges when a spatial dynamics perspective is applied to problems posed on $\mathbb{R}^{2}$, written in spherical coordinates $(r,\varphi)\in\mathbb{R}^{+}\times S^{n-1}$. This was first explored in the study of spiral wave patterns by Scheel in [18]. Later work by Sandstede and Scheel treated spiral waves in greater detail. They applied a dynamical systems perspective, and in particular the existence of exponential dichotomies, to rigorously characterize the spectral properties of spiral waves on both bounded and unbounded domains, the robustness of spiral waves, and the behavior of eigenfunctions to the linearization which indicate the spatial structure of patterns bifurcating from radially symmetric spiral waves [17].
@@
Recent work by Beck et. al. [2] has extended the use of spatial dynamics further to arbitrary Euclidean spatial domains, and in particular to elliptic PDEs posed on $\mathbb{R}^{n}$. The application of spatial dynamics in $\mathbb{R}^{n}$ looks much the same as in the setting of $\mathbb{R}^{2}$, relying again on spherical coordinates $(r,\varphi)\in\mathbb{R}^{+}\times S^{n-1}$. The elliptic PDE
$0=\Delta u-V(x)u,\qquad V:\mathbb{R}^{n}\to\mathbb{C}^{d\times d},\quad u\in\mathbb{R}^{n}\to\mathbb{C}^{d}$ (1)
is rewritten in spherical coordinates as
@@
$u_{rr}+\left(\frac{n-1}{r}\right)u_{r}+\frac{1}{r^{2}}\Delta_{S^{n-1}}u+V(r,\varphi)u=0,\qquad u(r,\cdot)\in L^{2}(S^{n-1},\mathbb{C}^{d})$
@@
posed on the Banach space $X=H^{1}(S^{n-1},\mathbb{C}^{d})\times L^{2}(S^{n-1},\mathbb{C}^{d})$ so that $\mathbf{u}=(u,v)\in X$.
@@
The results of Beck et. al. establish not only the method of spatial dynamics, but also the equivalence of weak PDE solutions (e.g. weak solutions to (1)) and solutions to the corresponding spatial evolutionary equation (e.g. solutions to (2)) in both $\mathbb{R}^{n}$ and general bounded Euclidean domains. These results lay the groundwork for adapting classical dynamical systems techniques to this newer spatial dynamics setting. In fact, in [1], Beck et. al. show the existence of an exponential dichotomy for $r\in(0,R]$ (the core field) and prove that the unstable space defined by this dichotomy corresponds to the boundary data of weak PDE solutions; there is a correspondence between “well-behaved” solutions of the spatial evolutionary system and weak PDE solutions.
Exponential dichotomies extend the notion of invariant stable and unstable linear subspaces to the nonautonomous and infinite dimensional settings [3, 9, 10, 11]. They characterize a hyperbolic splitting into an exponentially, asymptotically stable invariant evolution and an exponentially, asymptotically unstable invariant evolution. The incredible power of exponential dichotomies is that they allow us to characterize asymptotic behavior in systems where other tools for asysmptotic analysis may fail. For example, even for spatial evolutionary equations that are ill-posed, exponential dichotomies may exist and can then describe the behavior of bounded solutions. Exponential dichotomies have been used to understand spectral and Fredholm properties of operators (e.g. [3, 10, 15, 17]), to construct center manifold reductions (e.g. [6, 7, 8, 17, 18]), and to construct stable and unstable manifolds and therefore to also study homoclinic and heteroclinic solutions (e.g. [9, 12, 16]).
In this paper, we prove the existence of a far field dichotomy— an exponential dichotomy on the interval $r\in[R,\infty)$ for $R<\infty$— for the same spatial evolutionary equation (2). Existence proofs for exponential dichotomies typically rely heavily on the robustness of exponential dichotomies which was first established in an infinite dimensional setting by Peterhof, Sandstede, and Scheel in [11]. This property means that exponential dichotomies persist through small perturbation, so existence arguments are often reduced to finding the “right” decomposition of an operator into a principal part that is known to have an exponential dichotomy and a perturbation that is sufficiently small.
In [1], the primary challenge of proving the existence of the core-field dichotomy lay in the singularity at $r=0$ which was overcome by a reparameterization of $\tau=\ln r$ (see also [18]) and an exponential rescaling of the resulting system posed on $\tau\in\mathbb{R}^{-}$. In this new rescaled system, the asymptotic operator that emerges when $\tau$ is sent to -$\infty$ can be chosen to be the principal operator while the perturbation that remains decays exponentially in $\tau$.
@@
Applying robustness to prove the existence of a far-field exponential dichotomy is quite challenging due to the nonuniform unboundedness of the $r^{-2}\Delta_{S^{n-1}}$ term. Though this term vanishes as $r$ is sent to infinity and does not primarily control the hyperbolicity of the system, its unboundedness with respect to the $L^{2}$ norm means that it cannot simply be considered a “perturbation.” Instead, we must consider this Laplacian inside of the principal part and work within a function space where it is bounded but does not dominate the underlying hyperbolicity of the system. Because of the nonuniformity of the $r^{-2}\Delta_{S^{n-1}}$ term, we equip the underlying function space $X=H^{1}(S^{n-1})\times L^{2}(S^{n-1})$, on which (2) is posed, with an explicitly $r$-dependent norm, taking inspiration from the approach of Sandstede and Scheel in the spiral waves context [18, 17].
@@
Observing that $\mathcal{A}(r)$ is both closed and densely defined in $X$ with domain $X^{1}=H^{2}(S^{n-1})\times H^{1}(S^{n-1})$, we adapt the natural norms for $X$ and $X^{1}$ which are given by
$\|\mathbf{u}(r,\cdot)\|_{X}^{2}$ $=\|u(r,\cdot)\|_{H^{1}}^{2}+\|v(r,\cdot)\|_{L^{2}}^{2}$
$\|\mathbf{u}(r,\cdot)\|_{X^{1}}^{2}$ $=\|u(r,\cdot)\|_{H^{2}}^{2}+\|v(r,\cdot)\|_{H^{1}}^{2},$ (3)
to $r$-dependent norms by weighting the $H^{1}$ and $H^{2}$ norms, respectively:
$\|\mathbf{u}(r,\cdot)\|_{X_{r}}^{2}$ $:=\frac{1}{r^{2}}\|u(r,\cdot)\|_{H^{1}}^{2}+\|u(r,\cdot)\|_{L^{2}}^{2}+\|v(r,\cdot)\|_{L^{2}}^{2}$
$\|\mathbf{u}(r,\cdot)\|_{X^{1}_{r}}^{2}$ $:=\frac{1}{r^{2}}\|u(r,\cdot)\|_{H^{2}}^{2}+\|u(r,\cdot)\|_{H^{1}}^{2}+\|v(r,\cdot)\|_{H^{1}}^{2}.$ (4)
@@
In the $r$-dependent function space $X_{r}$, the dynamics of the asymptotic hyperbolic operator dominates norm estimates as $r$ goes to infinity while the influence of the spherical Laplacian remains bounded under the weighted $H^{1}$ norm. This choice is the key trick to applying robustness in this setting and allows us to prove our result.
@@
Overview of Paper. In section 2, we present this framework for the existence of a far-field dichotomy for (2) in greater detail. Later, in Section 3, we prove the Lemmas and existence results stated in Section 2. Finally, in Section 4, we discuss possible applications of our results, particularly with respect to the existence and construction of localized spatial patterns in $\mathbb{R}^{n}$.
@@
Our approach to proving the existence of exponential dichotomies relies heavily on the property of robustness, the persistence of exponential dichotomies through sufficiently small perturbations. Therefore, to prove that (2) has an exponential dichotomy on an appropriately chosen function space, we decompose it into a principal part that we can directly prove has an exponential dichotomy and a perturbation that we can bound appropriately.
Often the asymptotic operator, $\mathcal{A}_{\infty}=\lim_{r\to\infty}\mathcal{A}(r)$, is a suitable choice for the principal part of a robustness argument. Finding a principal operator is not so simple in this case, however. Assuming that $V(r,\varphi)$ decays to a constant potential $V_{\infty}\in\mathbb{C}^{d\times d}$ as $r$ goes to infinity, the asymptotic operator corresponding to (2) is given by
$\mathcal{A}_{\infty}=\begin{pmatrix}0&1\\ V_{\infty}&0\end{pmatrix}$ (5)
and is associated with the asymptotic equation:
$\mathbf{u}_{r}=\mathcal{A}_{\infty}\mathbf{u}.$ (6)
If we were to take the above $\mathcal{A}_{\infty}$ as the principal operator, we would be left with a remainder that is unbounded relative to $\mathcal{A}_{\infty}$ under the natural product norm on $X$ due to the spherical Laplacian term. Since the spherical Laplacian term can therefore not be left as a perturbation, we must include it in our principal operator and instead choose a norm that bounds the Laplacian when $r$ is small and highlights the effects of $V_{\infty}$ when $r$ is large. This is accomplished by the norms defined in (4). Notice that this choice of an $r$-dependent norm effectively creates a family of function spaces $X_{r}$ where the subscript denotes which value of $r$ is to be used in the norm. [^1] Recall that $X=H^{1}(S^{n-1},\mathbb{C}^{d})\times L^{2}(S^{n-1},C^{d})$. Consequently, a function $u:S^{n-1}\to\mathbb{C}^{d}$ may live in any $X_{r}$, however we will typically consider $u:[0,\infty)\times S^{n-1}\to\mathbb{C}^{d}$ evaluated at a fixed value of $r$ as living in the corresponding $X_{r}$ space; for example, $u(r,\cdot)\in X_{r}$ and $u(\rho,\cdot)\in X_{\rho}$.
With this new weighted family of function spaces, $X_{r}$, we can prove the existence of an exponential dichotomy for a principal part containing the leading order derivatives. This principal part is given by
@@
Before further discussing the existence proof, we must first define what is meant by an exponential dichotomy on a family of functions spaces such as $X_{r}$. To provide a solid foundation for this definition, we first state the standard definition of an exponential dichotomy on a general, fixed, Banach space $Y$.

### v2

# Exponential Dichotomies for Elliptic Equations on Multidimensional Domains

@@ Abstract
The existence of exponential dichotomies has been well-established as a powerful tool to study existence, stability, and bifurcations of coherent structures. Currently, the application of exponential dichotomies to elliptic problems posed on multi-dimensional domains is predominately limited to the context of cylindrical spatial domains. Recent work in [4] has shown how to extend the method of spatial dynamics, in which one views a spatial variable as a time-like evolutionary variable, to general multi-dimensional spatial domains. In this paper, we show that exponential dichotomies exist for a class of spatial dynamical systems arising in this more general setting, thus allowing for their use in future analyses of coherent structures.
@@ 1 Introduction
A current challenge in the study of pattern formation is the intractability of analytic dynamical systems approaches for pattern-forming partial differential equations in higher spatial dimensions. Dynamical systems techniques, for example center manifold reductions, geometric singular perturbation theory, and exponential dichotomies, have been applied to study the existence, stability, and bifurcations of coherent structures in a variety of settings posed on one-dimensional spatial domains, but their use in higher dimensions still remains limited to certain contexts. It is the aim of this work to broaden the class of higher-dimensional settings where such analyses can take place by furthering the theory of exponential dichotomies.
@@
Spatial dynamics for one-dimensional spatial domains. To study elliptic equations from a dynamical systems perspective, we employ the technique of spatial dynamics. In the simplest context, spatial dynamics is the trick of rewriting a second order differential equation into a system of first order equations. For example, the equation $u_{xx}+F(u)=0$, where $u(x)\in\mathbb{C}^{d}$ for each fixed $x\in\mathbb{R}$, becomes the system
$u_{x}$ $=v$
$v_{x}$ $=-F(u).$
where $(u(x),v(x))\in\mathbb{C}^{d}\times\mathbb{C}^{d}$. This same procedure provides a method for studying the stationary and traveling-wave problems associated with PDEs posed on $\mathbb{R}$. For example, a traveling wave ansatz $u(\xi)=u(x+ct)$ applied to the reaction-diffusion equation $u_{t}=u_{xx}+F(u)$ yields $cu_{\xi}=u_{\xi\xi}+F(u)$ which can be written as a first order system of ODEs as
$\begin{split}u_{\xi}&=v\\ v_{\xi}&=cv-F(u),\end{split}$ (1)
where $(u(\xi),v(\xi))\in\mathbb{C}^{d}\times\mathbb{C}^{d}$.
Equation (1) is the classical setting in which ODE tools can be explicitly applied to study pattern formation. For example, bifurcation theory, center manifold reduction, and geometric singular perturbation theory are traditional ODE tools which have been extensively applied in this setting to understand the emergence of non-uniform stationary spatial patterns on $\mathbb{R}$; see [8] for a helpful introduction. Of particular interest in this paper is the tool of exponential dichotomies, which were originally developed to extend the notion of invariant stable and unstable linear subspaces to the non-autonomous setting [7]. Because they characterize forward and backward exponential decay of linear (non-autonomous) systems, exponential dichotomies can play a significant role in constructing and analyzing bounded solutions that decay to known backgrounds states in the tails. Notably, exponential dichotomies can be used to construct bounded perturbations of pulses and fronts, which appear as homoclinic and heteroclinic solutions of (1), as in [16]. Other work on exponential dichotomies has shown their connection with the Fredholm properties of related differential operators; they therefore also play a significant role in understanding the temporal stability properties of stationary and traveling wave problems (e.g. [18]).
Spatial dynamics for cylindrical domains. Motivated by the success of dynamical systems tools for studying PDEs with one spatial dimension, [13] extended the notion of spatial dynamics to PDEs posed on cylindrical spatial domains of the form $\mathbb{R}\times\Omega\subset\mathbb{R}^{n}$ where $\Omega$ is a compact subset of $\mathbb{R}^{n-1}$. In this channel setting, stationary or traveling-wave solutions can be studied as solutions to a spatial evolutionary equation which describes how functions on the compact cross-section $\Omega$ evolve in the unbounded variable $x\in\mathbb{R}$. Consequently, spatial dynamics on the channel becomes an infinite-dimensional evolutionary equation rather than a finite-dimensional ODE in the case of one-dimensional spatial domains. For example, stationary solutions of $u_{t}=\Delta u+F(u)$ posed on a general $n$-dimensional channel $\mathbb{R}\times\Omega\subset\mathbb{R}^{n}$, will satisfy
$u_{xx}+\Delta_{\Omega}u+F(u)=0.$ (2)
By writing this equation as a first order system in the unbounded variable $x$, we can study the stationary problem as an evolutionary equation
$\begin{split}u_{x}&=v\\ v_{x}&=-\Delta_{\Omega}u-F(u)\end{split}$ (3)
posed on a Banach space, so that for each fixed $x\in\mathbb{R}$, $(u(x),v(x))\in H^{1}(\Omega,\mathbb{C}^{d})\times L^{2}(\Omega,\mathbb{C}^{d})$.
This approach yields linearized systems with compact resolvent, allowing for the application of center manifold techniques to reduce the infinite-dimensional dynamics of (3) to a finite-dimensional manifold governing small bounded solutions in both autonomous [13] and non-autonomous systems [14]. A similar finite-dimensional reduction by [15] extended these techniques to encompass not just small amplitude solutions but also larger bounded solutions. Front interfaces between these larger bounded solutions were also studied by [15] via an extension of the heteroclinic front construction techniques used in the analogous finite-dimensional setting.
Further extension of classic dynamical systems techniques to spatial dynamics on the channel proved challenging due to the ill-posedness of (3). The existence of well-behaved solutions in either forward or backward evolution is not guaranteed for these spatial evolutionary equations since the linear operator $\left(\begin{smallmatrix}0&1\\ -\Delta_{\Omega}&0\end{smallmatrix}\right)$ has an unbounded point spectrum with real part going to both positive and negative infinity. The theory of exponential dichotomies was extended to this setting by [17], and is a primary tool to overcome the challenge of ill-posedness. Just as in the finite dimensional setting, exponential dichotomies describe how to split the underlying space into invariant families of stable and unstable subspaces. Consequently, even for ill-posed evolutionary equations, exponential dichotomies describe where well-behaved solutions exist and decay exponentially in the forward direction (the stable family of subspaces) and where solutions exist and decay exponentially in the backward direction (the unstable family of subspaces). Furthermore, [17] showed how exponential dichotomies for infinite-dimensional problems can be combined with Lyapunov-Schmidt techniques in the style of [16] to construct bounded solutions of nonlinear equations that remain close to a pulse solution.
The primary reason why spatial dynamics works for studying elliptic PDEs on cylindrical domains is because the resulting linearized operator has compact resolvent. In the setting of (3) this is due to the fact that the cross-sectional spatial domain $\Omega$ is a compact subset of $\mathbb{R}^{n-1}$, but it can also be achieved in other ways. For example, a linearized operator with compact resolvent is obtained when studying time-periodic solutions of a reaction-diffusion equation posed on $\mathbb{R}$ by considering how functions in $L^{2}_{\mathrm{per}}([0,T],\mathbb{C}^{d})$ evolve in the unbounded spatial variable $x\in\mathbb{R}$ (e.g. [19, 20, 21]). A similar effect is also achieved when studying traveling fronts between spatially periodic background states. In this case, by treating the direction of propagation as the evolutionary variable, it is the imposition of periodicity in the cross-sectional space that ensures the linearized operator has compact resolvent. For example, [10] studied one-dimensional traveling fronts in which periodic patterns invade the trivial state, and [9] studied planar traveling fonts in which stable hexagonal patterns invade various unstable patterns.
Spatial dynamics on more general multidimensional domains. Spatial dynamics was extended to more general two-dimensional patterns, including target [24] and spiral waves [23], by decomposing $\mathbb{R}^{2}$ into spherical coordinates $(r,\varphi)\in\mathbb{R}^{+}\times S^{1}$ with the radial variable $r$ playing the role of the unbounded evolutionary variable. For example, radially symmetric stationary solutions of $u_{t}=\Delta u+F(u)$ satisfy $\Delta u+F(u)=0$, which when written in spherical coordinates (and imposing $u(r,\varphi)\equiv u(r)$) becomes
$\quad u_{rr}+\left(\frac{n-1}{r}\right)u_{r}+F(u)=0.$ (4)
This can then be written as a first order system in the variable $r$ as
$u_{r}$ $=v$
$v_{r}$ $=-\left(\frac{n-1}{r}\right)v-F(u),$
where for each fixed $r\in\mathbb{R}^{+}$, $(u(r),v(r))\in\mathbb{C}^{d}\times\mathbb{C}^{d}$. The results of [24] overcome both the non-autonomous dependence on $r$ and the singularity at $r=0$, developing a systematic bifurcation theory including extensions of center manifold reduction and normal form theory.
@@
A subsequent study of spiral waves from a spatial dynamics perspective was given by [22] in which the existence of exponential dichotomies were proven and applied to rigorously understand the spectral properties and robustness of spiral waves on both bounded and unbounded domains. In addition, [22] used this perspective to characterize the behavior of eigenfunctions to the linearization which indicate the spatial structure of patterns bifurcating from radially symmetric spiral waves.
@@
Recent work, [4], has extended the use of spatial dynamics further to arbitrary Euclidean spatial domains including, of course, $\mathbb{R}^{n}$. The application of spatial dynamics in $\mathbb{R}^{n}$ looks much the same as in the setting of $\mathbb{R}^{2}$, relying again on spherical coordinates $(r,\varphi)\in\mathbb{R}^{+}\times S^{n-1}$. The elliptic equation $0=\Delta u-V(x)u$ with $V:\mathbb{R}^{n}\to\mathbb{C}^{d\times d}$ and $u(x)\in\mathbb{C}^{d}$ for $x\in\mathbb{R}^{n}$ is rewritten in spherical coordinates as
@@
$u_{rr}+\left(\frac{n-1}{r}\right)u_{r}+\frac{1}{r^{2}}\Delta_{S^{n-1}}u-V(r,\varphi)u=0,\qquad u(r,\cdot)\in L^{2}(S^{n-1},\mathbb{C}^{d})$ (5)
@@
posed on the Banach space $X$ defined as
$X=H^{1}(S^{n-1},\mathbb{C}^{d})\times L^{2}(S^{n-1},\mathbb{C}^{d}),\qquad\big\|\mathbf{u}(r,\cdot)\big\|_{X}^{2}=\big\|u(r,\cdot)\big\|_{H^{1}}^{2}+\big\|v(r,\cdot)\big\|_{L^{2}}^{2}$ (7)
so that, for each $r\in\mathbb{R}^{+}$, $\mathbf{u}(r)=(u(r),v(r))\in X$.
@@
The results of [4] establish not only the method of spatial dynamics, but also the equivalence of weak PDE solutions and solutions to the corresponding spatial evolutionary equation in both $\mathbb{R}^{n}$ and general bounded Euclidean domains. These results lay the groundwork for adapting classical dynamical systems techniques to this newer spatial dynamics setting. In fact, [3] showed the existence of an exponential dichotomy for $r\in(0,R]$ (the core field), overcoming the singularity at $r=0$ in a relatively straightforward way using the reparameterization $\tau=\ln r$ and an exponential rescaling of the resulting system posed on $\tau\in\mathbb{R}^{-}$ (see also [23]).
Remark.
Though this discussion has predominately been framed in terms of the existence of solutions, questions of stability are also often addressed using the techniques described above. For example, one could study an eigenvalue problem for the linearization about a coherent structure in any of these settings by adding the linear term $\lambda u$ to Equations (2), (4), or (5).
Our contribution: Far-field exponential dichotomies for elliptic problems posed on $\mathbb{R}^{n}$. To extend the results of [3] to problems posed on the unbounded domain $\mathbb{R}^{n}$, we prove the existence of a far-field dichotomy—an exponential dichotomy on the interval $r\in[1,\infty)$—for the same spatial evolutionary equation (6). The basis of our proof relies on the robustness of exponential dichotomies, the property that exponential dichotomies persist through sufficiently small perturbations. While the standard robustness result in the cylindrical setting given by [17] can be applied in a relatively direct way for the core-field dichotomy proof in [3], our proof of the far-field dichotomy is complicated by the non-uniform unboundedness of the $r^{-2}\Delta_{S^{n-1}}$ term in (6).
@@
Though the $r^{-2}\Delta_{S^{n-1}}$ term vanishes as $r$ is sent to infinity and does not primarily control the hyperbolicity of the system, its unboundedness with respect to the $L^{2}$ norm means that it cannot simply be considered a perturbation. Instead, we must show that the principal component of (6), including this spherical Laplacian term, has an exponential dichotomy in an appropriately chosen function space. Taking inspiration from the robustness argument in [23] and [22], we choose to work in a function space with an explicitly $r$-dependent norm which complements the non-uniformity of the $r^{-2}\Delta_{S^{n-1}}$ term by ensuring that it remains bounded while not dominating the asymptotic hyperbolicity of the system.
@@
In particular, observing that $\mathcal{A}(r)$ is both closed and densely defined in the space $X$ with domain $X^{1}=H^{2}(S^{n-1})\times H^{1}(S^{n-1})$, we adapt the natural $X$ norm, as defined in (7), to an $r$-dependent norm by weighting the $H^{1}$ component:
$\begin{array}[]{c}X_{r}=H^{1}(S^{n-1},\mathbb{C}^{d})\times L^{2}(S^{n-1},\mathbb{C}^{d}),\\[10.00002pt] \big\|\mathbf{u}(r,\cdot)\big\|_{X_{r}}^{2}:=\frac{1}{r^{2}}\big\|u(r,\cdot)\big\|_{H^{1}}^{2}+\big\|u(r,\cdot)\big\|_{L^{2}}^{2}+\big\|v(r,\cdot)\big\|_{L^{2}}^{2}\end{array}$ (10)
@@
In this $r$-dependent function space $X_{r}$, the dynamics of the asymptotic hyperbolic operator dominate norm estimates as $r$ goes to infinity while the influence of the spherical Laplacian remains bounded under the weighted $H^{1}$ norm. This choice of norm turns out to be the right function space in which to show our system has an exponential dichotomy. A carefully constructed, norm-preserving transformation between $X_{r}$ and $L^{2}(S^{n-1})\times L^{2}(S^{n-1})$—presented in Lemma 2.3—is the key trick to applying robustness in this setting and allows us to prove our result.
It is worth noting that both [23] and [22] design similar $r$-weighted norms to establish the existence of exponential dichotomies in the spiral wave settings. Our choice of $r$-dependent norm differs from theirs in the middle unweighted $L^{2}$ norm of $u$ where they instead have an unweighted $H^{1/2}$ norm. This is due to an angular advection term that appears in the spiral waves context. Beyond the absence of an advection term in our setting, the main distinction between the proof of exponential dichotomies in the spiral waves setting and the proof presented here is the way that we decompose the system into a principal part and a perturbation. In [22], the principal component includes only the differential operators, treating the additional terms which are analogous to our asymptotic potential, $V_{\infty}=\lim_{r\to\infty}V(r,\varphi)$, as perturbations. The principal component in our analysis includes both the spherical Laplacian term and the asymptotic potential. As a result, our proof involves slightly more technical bounds to prove Lemma 2.3 (comparable to Lemma 5.6 in [22]), while the proof in [22] requires a more elaborate argument to bound the perturbation terms.
@@
Organizational overview of paper. In section 2, we present this framework for the existence of a far-field dichotomy for equation (6) in greater detail. Later, in Section 3, we prove the Lemmas and existence results stated in Section 2, with many of the technical details contained in the Appendix. Finally, in Section 4, we discuss possible applications of our results, particularly with respect to the existence and construction of localized spatial patterns in $\mathbb{R}^{n}$.
Acknowledgements. MB and AHH gratefully acknowledge NSF support under award number DMS-2205434. RG gratefully acknowledges NSF support under Award Numbers DMS-2006887 and DMS-2307650.
@@
Our approach to proving the existence of exponential dichotomies relies heavily on the property of robustness, the persistence of exponential dichotomies through sufficiently small perturbations. Therefore, to prove that (6) has an exponential dichotomy on an appropriately chosen function space, we decompose it into a principal part, $\tilde{\mathcal{A}}(r)$, that we can directly prove has an exponential dichotomy, and a perturbation, $\tilde{B}(r)$, that we can bound appropriately. This decomposition can be characterized as
[…]
where $\mathcal{A}_{\infty}=\lim_{r\to\infty}\mathcal{A}(r)$ is the asymptotic operator of $\mathcal{A}(r)$ under the assumption that $V(r,\varphi)$ decays to a constant potential $V_{\infty}\in\mathbb{C}^{d\times d}$ as $r$ goes to infinity.
Often the asymptotic operator is a suitable choice for the principal part of a robustness argument. Since spectral hyperbolicity is equivalent to the existence of exponential dichotomies in a constant coefficient system, it is straightforward to look for an exponential dichotomy of the asymptotic system:
$\mathbf{u}_{r}=\begin{pmatrix}0&1\\ V_{\infty}&0\end{pmatrix}\mathbf{u}=\mathcal{A}_{\infty}\mathbf{u}.$ (11)
In this context, however, it is not sufficient to consider $\mathcal{A}_{\infty}$ as the principal part, as we cannot directly apply a robustness argument to show that (6) has an exponential dichotomy when (11) does. In particular, if we were to take $\mathcal{A}_{\infty}$ as the principal operator, we would be left with a remainder that is unbounded relative to $\mathcal{A}_{\infty}$ under the natural product norm on $X$ due to the spherical Laplacian term. Since the spherical Laplacian term can therefore not be left as a perturbation, we must include it in our principal operator and instead choose a norm that both bounds the Laplacian when $r$ is small and highlights the effects of $V_{\infty}$ when $r$ is large. This is accomplished by the norms defined in (10).
With this new weighted family of function spaces, $X_{r}$ in (10), we can prove the existence of an exponential dichotomy for a principal part containing the leading order derivatives. This principal part is given by
@@
Before further discussing the existence proof, we must first define what is meant by an exponential dichotomy on the $r$-dependent space $X_{r}$.
Remark.
The choice of an $r$-dependent norm effectively creates a family of function spaces $X_{r}$ where the subscript denotes which value of $r$ is to be used in the norm. Consequently, a function $u:S^{n-1}\to\mathbb{C}^{d}$ may live in any $X_{r}$, however we will typically consider $u:[0,\infty)\times S^{n-1}\to\mathbb{C}^{d}$ evaluated at a fixed value of $r$ as living in the corresponding $X_{r}$ space; for example, $u(r,\cdot)\in X_{r}$ and $u(\rho,\cdot)\in X_{\rho}$. For this reason, we refer to $X_{r}$ both as a family of function spaces and as a single function space.